Your father is 50 years old and will retire in 10 years. He expects to live for 25 years after he retires, until he is 85. He wants a fixed retirement income that has the same purchasing power at the time he retires as $50,000 has today. (The real value of his retirement income will decline annually after he retires.) His retirement income will begin the day he retires, 10 years from today, at which time he will receive 24 additional annual payments. Annual inflation is expected to be 4%. He currently has $205,000 saved, and he expects to earn 7% annually on his savings. The data has been collected in the Microsoft Excel Online file below. Open the spreadsheet and perform the required analysis to answer the question below.

Open spreadsheet

How much must he save during each of the next 10 years (end-of-year deposits) to meet his retirement goal? Do not round your intermediate calculations. Round your answer to the nearest cent.

Excel Online template Saved File Home Insert Data Review View Tell me what you want to do Comments 二鹊 Sort&Filter s-% 9 %9 Formatting as Table Conditional Format Insert Delete Format Find &Select UndoClipboard B11 3 Retirement income today 4 Years to retirement 5 Years of retirement 6 Inflation rate 7 Savings 8 Rate of retum $50,000 10 25 4.00% $205,000 7.00% 0 Calculate value of savings in 10 years Formulas 11 Savings att 10 12 3 Calculate value of fixed retirement income in 10 years 14 Retirement income at t 10 15 Calculate value of 25 beginning-of-year retirement 16 payments at t 10: 17 Retirement payments att 10 18 19 20 Calculate net amount needed at Value of retirement payments 10 Sheet1 Help Improve Office
Retirement payments at t10 18 19 Calculate net amount needed att 10: 20 Value of retirement payments 21 Value of savings #NIA #N/A #N/A Net amount needed 23 24 Calculate annual savings needed for next 10 years 25 Annual savings needed for retirement 26 27 28 Sheet1 Help Improve Office
Answer #1
Value of savings in Years = 205000*1.07^10 = $   4,03,266.03
Value of fixed return income in 10 years = 50000*1.04^10 = $       74,012.21
Value of 25 beginning of year retirement income = 74012.21*1.07*(1.07^25-1)/(0.07*1.07^25) = $   9,22,882.97
Net amount needed:
Value of retirement payments $   9,22,882.97
Value of savings $   4,03,266.03
Net amount needed $   5,19,616.94
Annual savings needed for next 10 years = 519616.94*0.07/(1.07^10-1) = $       37,608.58